### v2

# Immunization against harmful fine-tuning attacks

Abstract
Large Language Models (LLMs) are often trained with safety guards intended to prevent harmful text generation. However, such safety training can be removed by fine-tuning the LLM on harmful datasets. While this emerging threat (harmful fine-tuning attacks) has been characterized by previous work, there is little understanding of how we should proceed in constructing and validating defenses against these attacks especially in the case where defenders would not have control of the fine-tuning process. We introduce a formal framework based on the training budget of an attacker which we call “Immunization” conditions. Using a formal characterisation of the harmful fine-tuning problem, we provide a thorough description of what a successful defense must comprise of and establish a set of guidelines on how rigorous defense research that gives us confidence should proceed.
Immunization against harmful fine-tuning attacks
Domenic Rosati1,5 Jan Wehner2 Kai Williams3 Łukasz Bartoszcze4 Hassan Sajjad1 Frank Rudzicz1,5 1Dalhousie University 2TU Delft 3Swarthmore College 4University of Warwick 5Vector Institute for Artificial Intelligence 1domenic.rosati@dal.ca
1 Motivation
An emerging research direction demonstrates that safety techniques for Large Language Models (LLMs) can easily be circumvented by fine-tuning them on harmful samples (Yang et al., 2023; Lermen et al., 2023; Qi et al., 2023) - Examples of this occurring in practice are provided in Appendix A. To ensure the safe release and deployment of LLMs, we should invest in defenses against these attacks. These defences are necessary due to the following:
Additionally, concerns about model liability and responsibility (Henderson et al., 2023a) might prioritize questions about how model developers restrict downstream fine-tuning beyond easily ignored licensing schemes.
Previous work has attempted to understand the emergence of harmful text generation (Wei et al., 2024) and defend against it (Henderson et al., 2023b; Zhou et al., 2023; Huang et al., 2024). However, this work lacks clear conceptual analysis of defense conditions such as what is the definition of a successful defense and what are the necessary conditions defenses must meet. Without this, it is unclear how we should validate proposed defenses. For example without careful experimental design future works might present effective defences that completely ruin model capability or do not generalize to unseen attacks.
In this paper, we ground the “Harmful Fine-Tuning Attack" (HFTA) threat model (see Figure 1) in a training budget framework that highlights the cost considerations of defense (Henderson et al., 2023b; Apruzzese et al., 2022). We propose a novel set of necessary conditions for defenses against these attacks we call “Immunization” conditions (§ 3). Our contribution is a formal illustration of what defenses against harmful fine-tuning attacks must look like: i.e. resistance to the process of supervised fine-tuning, stablity which retains general language modeling capability, generalizable across various attack datasets, and trainablity to maintain the utility of the LLM for learning harmless tasks. Using these conditions, we present clear guidelines on how to operationalize them to conduct future defense research § 4 ( empirical demonstration presented in Appendix C).
2 Threat Model for HFTAs
Goal
The goal of a HFTA is to use an LLM to cause harm or conduct illegal activities which would otherwise be prevented by safety training. Phishing provides a specific example (Hazell, 2023) but we refer readers to Weidinger et al. (2022) for a proper discussion of what constitutes harm. Our working definition for this threat model is that harm is defined by the defender as we will discuss more below. This may involve removing existing safety guards or further training on a harmful dataset to improve a harmful capability. Appendix B provides a literature review outlining previous threat models, attacks, and defense research. Figure 1 illustrates where harmful fine-tuning attacks fits within the landscape of other types of attacks: backdoor attacks implant a trigger at training-time, occur before model release, and require stealth (a trigger must be hidden such that a defended cannot find it). Adversarial attacks occur at inference-time only post-model release. In contrast to these harmful fine-tuning attacks occur post-model release that don’t require stealth and are training-time attacks.
Strategy
Attackers achieve harmful fine-tuning by using the training objective below, Equation 1, on a given LLM without harmful capabilities which we denote as , where indicates the number of optimization steps taken ( is the initial model) to find the parameters for model . The attacker will use a harmful dataset which includes prompts and target responses designed to elicit harmful behavior in . They then minimize the loss function by taking training steps up to their compute budget () (Henderson et al., 2023b). The outcome is the optimal parameters found at some training step that the attacker can use to engage in illegal and harmful activities that would have previously been refused. This we call harmful training - eq. 1 - where is found by:
Capabilities
We assume that the attacker has a limited compute budget to train a harmful model i.e. not enough to train an LLM from scratch. We also assume enough training data to achieve meaningful harmful performance. In the white box setting, the attacker can modify the architecture, loss functions for training, inference process, etc. since they have complete access to the trained model. This generally corresponds to the case of open source release or model theft. Black box settings assume the attacker can only access a fine-tuning API mediated by the defender. Here the attacker pays for training for example using input token-based pricing found in OpenAI’s pricing model 111https://openai.com/api/pricing/.
3 Immunization Conditions
Given the threat model above, we present conditions for successful defense that crystallize the work presented in Henderson et al. (2023b); Zhou et al. (2023); Huang et al. (2024). Appendix B explores fully how previous work fits within our framework. We say that a model is immunized (we indicate an immunized model as ) against harmful training (eq. 1) with respect to if it meets the following conditions:
Condition (1) - Resistance: Resistance measures the degree to which a defense can prevent training towards harmful ends. It follows from the requirement to prevent learning harmful behaviour and is the implicit goal of previous defense works.
For this condition, we first need to define a threshold which is the maximum acceptable harmfulness as estimated by some proxy metric measured by . This is a quantity set by the defender. For example, could be the result of a toxicity classifier as . We will say more about specifying harm and empirical attack success measures in § 4. For this we have:
Strong resistance means that eq. 1 never converges below a given loss whose value indicates a model has become harmful as measured by being generating text above the acceptable harmfulness threshold . This is for a given immunized model where is the training steps taken to find eq. 1 and is estimated as above. This condition provides defense against an unlimited budget of training steps:
Weak resistance does not require that the immunized model is always below a specified threshold . Instead, training takes at least train steps to pass the threshold . Here the defender wants to maximize the number of training steps in eq. 1 required to pass the threshold. In practical terms, the defender is ensuring the defense is too expensive for the attackers budget:
Resistance applies to both actively trying to learn a brand new harmful task as well as circumventing or “unlearning” safety guards. Note that resistance presupposes the model already performs poorly on these domains (as measured by which we do not specify as a separate condition itself.
Condition (2) - Stability: The stability condition states that we want to be able to continue to use the immunized model at the same or similar level of performance on harmless tasks as the unimmunized model . This ensures that the model remains usable. Using the proxy LLM capability measure we have:
where is some reference dataset on which we want to ensure stability. This condition is drawn from the empirical work of Zhou et al. (2023) which presented a suite of proxy metrics and different types of for measuring degradation. As we will discuss § 4, stability should also imply that models are not made more unsafe for example by inference-time jail breaking.
Condition (3) - Generalization: We assume the defender does not have access to the samples used by the attacker. Therefore we say that given disjoint subsets the defense procedure producing based only on should be resistant to HFTAs using . In-domain generalization requires that these subsets are drawn from the same domain. For example, a defense might be performed using examples of toxic content while the attack would be performed using non-overlapping samples of toxic content. This exemplifies the case where defender knows the domain the attacker will draw samples from. However this might not be strong enough for the defender who wants a model that can be robust across many types of harms that might only have small distributional overlap. Cross-domain generalization requires that the subsets be drawn from different domains for example immunization applied using samples from toxic content generation defending against an HFTA performed using harmful QA. We provide recommendations § 4 below on how defenses should demonstrate generalization.
Condition (4) - Trainability: This condition simply says that we are able to improve performance during supervised fine-tuning on an arbitrarily chosen harmless dataset within a similar number of train steps for the immunized model as , the unimmunized model and was shown in Zhou et al. (2023); Henderson et al. (2023b). In practice, we can define our tolerance of the difference between these two as some small .
This is an optional condition since defenses without trainability would still protect against HFTAs. We highlight trainability because of the high social utility of model training for commercial and research applications. Without trainability, there would be high social social pressure to non-maliciously undo defenses simply due to wanting to perform harmless training.
4 Guidelines For Future Research
Establishing resistance
First we recommend constructing methods that have theoretical guarantees like bounds on the number of samples necessary to undo a defense. Studies in transfer learning have characterized the difficulty of training towards a target task given a network trained on a source task. Both Achille et al. (2019) and Ben-David et al. (2010) provide theoretical frameworks that we encourage researchers to use. For example, Achille et al. (2019) provides a theoretical explanation of resistance by locating the parameters of an immunized model as a point in the dynamics of the loss landscape that minimize the transition probability of traversing a training trajectory in the loss landscape where the transition probability is a function of the curvature of the loss landscape. We encourage researchers to derive loss functions that provably minimize this transition probability. Researchers can develop strong resistance using cryptographic methods such as those found in Mukherjee and Chakraborty (2023); Alam et al. (2020) where the weights themselves cannot be modified without an authorization mechanism.
Harmful dataset selection and creation
Empirical evaluation of the resistance condition would begin with datasets used to construct realistic simulations of HFTAs. Currently, the two main types of datasets with validated harm measurement constructs consist of harmful question answering (Wang et al., 2024b; Ji et al., 2023) and toxic text generation (Gehman et al., 2020). Robust defense evaluations should investigate many more types of harms for example those outlined in Kapoor et al. (2024). Other types of harmful datasets such as fraud are more difficult to come by and are potentially high risk to construct since they could inadvertently assist attackers. Mitigation techniques for dataset construction include keeping datasets private, developing toy harmless domains that only simulate harms, strong licensing, and focusing on datasets that are already publicly available.
Measuring resistance
Since we focus on supervised fine-tuning attacks we recommend two initial dimensions of attack strength: learning rate and number of samples (or epochs) used. Other dimensions of attack could include batch size, the type of optimizer used and various other hyperparameters but they are harder to specify and future research will need to develop stronger attacks to ensure robust defenses. We advocate that simulated HFTAs use as many samples (or epochs) as possible across a large range of learning rates, a large dataset, and many epochs. We encourage future studies to comprehensively sweep learning rate from a rate that is too low for any learning to a learning rate that is high enough to ruin stability.
Attack success is a key measure of resistance and should be evaluated in the following way. First, as above, attack success consists of using a proxy evaluation function for harmfulness to measure whether we have surpassed a defender set threshold . Our recommendation is that we use domain specific measures validated by previous community efforts over generic measures of harm (Inan et al., 2023). For setting a threshold of acceptable defense we could either: (i) set the threshold of acceptable behavior as the harmfulness of the original undefended safety guarded model before a HFTA or (ii) A weaker threshold could be used where the defended model should simply be strictly less harmful than the base model after performing a successful HFTA. We don’t think that (ii) should be considered an actual defense but for challenging datasets it helps us construct an initial set of research goals. In practice, we expect defenses to compete empirically over their “degree of resistance” by minimizing the overall harmfulness measure as much as possible.
Ensuring Stability
For stability, we encourage measuring performance differences between the base model and the defended models on standard LLM capability benchmarks such as using several standard datasets (like MMLU Hendrycks et al., 2021) with Eleuther’s LM Harness. 222https://github.com/EleutherAI/lm-evaluation-harness While perplexity on a language modeling dataset such as WikiText (Merity et al., 2016) could be useful, it is a more indirect proxy of how the language model will behave. Additional requirements of stability should include that the model not be less safe across dimensions like inference-time adversarial attacks (for example those measured by (Mazeika et al., 2024) or general trustworthiness (Wang et al., 2024a) and bias (Esiobu et al., 2023).
Adaptive Attacks
All defences should attempt to be fully transparent about the conditions under which the defences can be broken and attempt to formulate adaptive attacks, i.e. attacks which could be successful if the attacker knew the defence method used. For example methods that rely on guarantees from the assumption the attacker uses gradient methods might employ Newton’s method to demonstrate an attack.
Encouraging Generalization
Generalization is particularly important to measure because in practice it is very unlikely that the defender will have access to the same samples as the attacker. To simulate this, the defender can consider the following: For in domain generalization, the defense and the attack should be performed on disjoint subsets of the same domain. Typically this will take the form of exploring sample efficiency: how small of an immunization dataset can we construct such that we protect against an HFTA in that domain? For cross domain generalization, there is an open question about what can be realistically expected from immunization methods. One expectation we might have is that our immunization methods should robustly defend across various sub-domains within a domain like harmful question answering. For example, with BeaverTails (Ji et al., 2023), a given harmful subset such as animal abuse or criminal activity could be left out as the attack dataset and we could evaluate whether we are able to defend against this left out subdomain. More challenging types of cross-domain generalization should also be evaluated such as if the defense was performed for toxic content generation and the attack was performed using harmful question answering.
Evaluating Trainability
To demonstrate trainability, we must select tasks that LLMs do not perform well on without additional training (i.e. we observe a large increase in performance after training). Like stability, we should select from benchmarks of natural language generation tasks with well-validated measures such as the GEM benchmark (Gehrmann et al., 2021). The GEM benchmark is recommended because we can construct text-to-data or structured data generation tasks that LLMs are unlikely to perform well on zero-shot.
Appendix C provides an empirical demonstration of some of these recommendations.
5 Conclusion
In this paper, we clarify an emerging threat model that identifies the risk of harmful fine-tuning in LLMs and quantifies it using an imagined attackers training budget. Recent work has shown that there might be ways in which we prevent these attacks. To assist with defense development, we proposed a formal set of “immunization” conditions that allow us to validate whether we have defended against harmful fine-tuning. We leverage these conditions to provide a concrete set of guidelines that future work could follow to develop robust evaluation of proposed defenses.
6 Limitations
Our setting currently only applies to supervised fine-tuning attacks. By its nature, this framework cannot consider settings where an attacker has the budget to train LLMs from scratch which would be better tackled by compute governance solutions (see Sastry et al., 2024). We assume that training high-capability LLMs from scratch will remain a significant cost factor that will exclude this as an option for most attackers. Other attack settings that resemble fine-tuning such as with model editing Li et al. (2024) are not considered in our framework but depending on the method could be considered as a future extension. We also don’t consider attack settings based on reinforcement learning which future defenses should consider due to the popularity of methods like DPO Rafailov et al. (2023) and evidence of using these for constructing attacks Yi et al. (2024).
One limitation of this research direction generally, is that it requires consensus about what constitutes harm and the collection of datasets exemplifying harmfulness. We acknowledge that defining harm is a contentious issue that might privilege harm against some groups at the expense of others and is a general issue endemic to LLM safety research (see discussion here Kirk et al., 2023).
While developing immunized models could provide much safer open weight release scenarios, the collection of datasets to immunize these models could present a dual-use risk if they are shared publicly especially if they are used to demonstrate successful attacks on undefended models with openly available code which bad actors can subsequently copy and use. Because of this we caution others to ensure safe and responsible dissemination of research results and artifacts.
We also highlight that this setting is a version of the more general domain authorization research program (Wang et al., 2021, 2023) where we only consider the harmfulness domain and only at training-time. Future work could consider extending this framework to domains that are not explicitly harmful but nonetheless are high risk such as health, legal, financial and military applications and therefore we might want to prevent training towards.
Finally, we clarify that “immunization” assumes that models are already made safe at inference-time and that we are attempting to prevent undoing that safety. This is not currently the case as there are many open safety challenges with LLMs (Anwar et al., 2024) that need to be addressed in tandem with this work. Our work also doesn’t speak to additional challenges in LLM security like jail-breaking which also needs to be developed along side HFTA defenses. Because of large attack surface of LLMs, we need to be sure we work together as a community in developing “defense-in-depth” solutions.
Appendix A Analysis of Concerning Models in the Wild
This section shows multiple cases where openly available LLMs hosten on huggingface have been intentionally trained for purposes that could be considered harmful or offensive, categorises types and methods of these attacks and showcases that popular training techniques can indeed be used to train harmful models. Our working definition for harmfulness for this section is simply that the content possibly generated by these models would be rejected or not produced by a conventional safety-aligned model. However, we acknowledge that different communities define harmfulness differently and this does not mean that these models do not have legitimate use cases, nor that these are necessarily against the original model license or Huggingface’s policies.
We conducted a non-exhaustive search of harmful text generation models on Huggingface 333Searches performed on February 19th 2024 on https://huggingface.co/models?pipeline_tag=text-generation&sort=trending. We search for LLMs with the keywords "uncensored", "unfiltered", "lewd", "NSFW", "evil" and "toxic". Instead of showing all models, we provide canonical examples of different types of potential harm and training methods in Table 1.
The most popular type of potential harm is uncensored models, which resulted in 267 search results. This results in models like WizardLM-13B-Uncensored that follow instructions but do not have safety guards built-in and thus don’t refuse to generate harmful content. This enables more flexible use cases like fictional writing and role-play but could be used to generate illegal content. Furthermore, we found 43 models trained to be toxic (search term: "toxic) and 21 trained to be evil (search term: "evil"). This is often done for purposes of role-playing or for research purposes (e.g. Pivot-0.1-evil-a). Lastly, there are 29 "NSFW" and 21 "lewd" models, often to be used for erotic role-playing. While these erotic models might not be considered harmful by some, they might generate content against the license of open-source models 444for example section (1) in https://llama.meta.com/llama-downloads/. We highlight Huggingface’s set of ethical guidelines 555https://huggingface.co/content-guidelines for hosting models, which has led them to ban models like gpt-4chan explicitly trained for harmful ends but doesn’t explicitly forbid most of the cases we highlight. We did not find models explicitly trained for fraud or other criminal activities.
A common method to train toxic or erotic models is to fine-tune an LLM on toxic (toxic-dpo-v0.1, ToxicQAtextFiltered) or erotic (instruct_nsfw_cn) datasets via DPO Rafailov et al. (2023), LoRA Hu et al. (2021) or Supervised Fine-Tuning (SFT). Other approaches train a toxic model by reversing the DPO objective while training on alignment datasets. Another common method is model merging techniques, where two or more models and their capabilities are combined into one model Ilharco et al. (2023). By merging a generally capable model like Llama with a model trained on lewd content, one can combine their capabilities to achieve an erotic role-playing LLM. Lastly, uncensored models can be attained by training a pre-trained LLM on variants of instruction-tuning datasets like Alpaca Wizardlm_alpaca_evol_instruct_70k or Vicuna Sharegpt_vicuna_unfiltered where data-points meant to align the LLMs were removed, resulting in models like WizardLM-13B-Uncensored. These resemble the begning fine-tuning attack presented in Qi et al. (2023) indicating that this type of attack was known to the community well before this work was presented.
It is interesting to note that multiple potentially harmful models were trained by methods like DPO or LoRA. This highlights the fact that these methods can be used symmetrically to align models or for harmful purposes and should also investigated for defense research in addition to harmful supervised fine tuning. We also point out that this is a case in practice of the dual use risk of alignment techniques since advancements like DPO are used to construct harmful models.
We find multiple cases of HFTA consistent with our threat model. For example, MLewd-L2-Chat-13B uses the Xwin-LM model which was trained via SFT and RLHF and phi-2-uncensored is a fine-tune of Phi-2 which was only trained on safe and educational websites. Furthermore, the amount of models which meet our search results indicates that there is a strong demand for harmful models.
| Name | Training Method | Dataset | Base Model | Safety Methods on Base Model |
| WizardLM-13B-Uncensored | Instruction Tuning | WizardLM alpaca evol instruct 70k unfiltered | Llama 2 | Pre-trained |
| phi-2-uncensored | DPO | toxic-dpo-v0.1 | Phi-2 | Only trained with safe content |
| toxicqa-Llama2-13B | LoRA | ToxicQA text Filtered | LLama2 | None |
| ToxicHermes-2.5-Mistral-7B | DPO | toxic-dpo-v0.1 | Mistral-7B-v0.1 | None |
| PiVoT-0.1-Evil-a | Reverse DPO | hh-rlhf, ko_wikidata_QA, OpenOrca-KO | Mistral 7B | None |
| MLewd-L2-Chat-13B | Model Merging | _ | Xwin-LM | SFT, RLHF |
| NSFW_13B_sft | SFT | instruct nsfw cn | Baichuan-13B-Base | None |
Appendix B Survey of Harmful Fine-tuning Attacks, Defenses, and Analysis
This section provides a literature review on the attack, defense, and analysis research produced thus far as related to harmful fine-tuning attacks tying back previous work to ours where possible.
B.1 Attacks
Training-time attacks specifically for removing safety guards through fine-tuning have previously been identified (Dong et al., 2024) and have been grouped with several other distinct types of training-time attacks like data poisoning and backdoor attacks. Due to the nature of our threat model not requiring stealthiness (the attacker does not hide their intentions during the attack), we don’t review previous literature on data posioning (Shu et al., 2023; Wan et al., 2023) and backdoors (Hubinger et al., 2024; Xu et al., 2023; Li et al., 2024; Bagdasaryan and Shmatikov, 2022; Rando and Tramèr, 2023; Cao et al., 2023) for training LLMs for harmful purposes. To provide clarity for those unfamiliar with the LLM security landscape, training-time attacks are attacks that occur during the training of an LLM. These are not related to inference-time attacks like jail breaks (Zou et al., 2023) and prompt injection (Greshake et al., 2023) which are done during the usage of a model after training and represent a different but equally important attack surface.
The first works we are aware of that show the ability to remove safety guards and train safety aligned LLMs towards harmful ends emerged in late 2023. Works such as Yang et al. (2023) presented a “shadow alignment” attack which uses supervised fine-tuning on only 100 question answer pairs. These are generated based on questions forbidden under OpenAI’s usage policy to demonstrate circumvention attacks across 8 models which transfer cross-lingually. Qi et al. (2023) perform an “Identity shifting attack” by constructing dialogue samples where the assistant is an absolutely obedient agent. Concerningly they also show that benign instruction following dialogues is able to circumvent the safety guards in both GPT-3.5-Turbo through the fine-tuning API as well as Llama-2-7b-chat. Lermen et al. (2023); Gade et al. (2023) construct a dataset, RefusalBench, to measure the effectiveness of circumvention attacks and show similarly that refusal rates drop to near 0 after spending less than $200 on training a series of safety guarded Llama2 models. They show the circumvention attack works using LoRA with the rest of the parameters frozen. Bhardwaj and Poria (2023) call the same attacks parametric red-teaming or unalignment and corroborate the results of earlier attackers of being able to circumvent a wide variety of publicly available safety guarded models as well as GPT-3.5-turbo with a very small number of samples. Importantly, they point out that these attacks are much more efficient than current jailbreak discovery and red teaming methods which many take many months of manual work or are more computationally expensive to find. Zhan et al. (2023); Pelrine et al. (2023) both show that for the new GPT-4 fine-tuning API, safety guards are similarly easy to remove in as little as 15 harmful or 100 harmless instruction following samples. Finally, Yi et al. (2024) showed that in addition to supervised fine-tuning, safety guards can be easily removed by simply reversing the preferred and dispreferred samples for preference learning with Direct Policy Optimization.
B.2 Defenses
Defense literature against this threat model is still emerging with two general defense settings: white- and black-box access. First are interventions design to work with white box threat models for which only our demonstration appendix C and Henderson et al. (2023b) provide proofs of concept thus far (Although Henderson et al. (2023b)’s work is directed to the classification and not the natural language generation (NLG) setting, making our sample demonstration the first white box defense demonstration for NLG). We note that we prefer to speak of immunization rather than self-destruction as a defense framing because (A) some defenders may want to preserve the ability to train towards harmless ends and (B) we may require the model to continue to work after harmful training and not self-destruct, i.e. result in a model that is not capable of any task after harmful training.
For black box settings, training is mediated by an API so the defender has control over the entire training pipeline. Therefore they have access a wider variety of interventions including much cheaper data filtration and post-deployment monitoring solutions. Zhou et al. (2023) proposes security vectors, where they train a set of removable parameters using LoRA to minimize loss on a given harmful dataset. The intuition is that if the loss on harmful samples is already very low, this will prevent meaningful gradient updates of the LLM during training. The authors provide a small demonstration that security vectors can generalize in-domain to preventing training on harmful samples without significantly reducing capabilities. Similar to Zhou et al. (2023), Bhardwaj et al. (2024) demonstrates the ability to use simple vector arithmetic with a safety vector estimated by analyzing the shift from the aligned mode to unaligned mode to realign a model that has been trained to be harmful. Huang et al. (2024) observed that supervised fine-tuning on harmful samples resulted in embeddings drifting much farther from the original aligned embeddings than training on harmless samples did. Based on this observation, they formulate a bilevel adversarial perturbation loss, “Vaccination”, where the inner loss finds the perturbation that would cause the most embedding drift from the original aligned model embeddings and the outer loss adds this term for minimization to regular cross-entropy language modeling loss. Finally, Zhao et al. (2023) leverage the fact that both safe guards and harmful generations are forgotten by the model after a certain amount of training on the opposite dataset. Using this observation, they construct a filter to optimize selection from a dataset that may contain harmful samples in order to prevent forgetting safety guards
B.3 Analysis
Finally, we present a number of analysis techniques used to localize and understand the harmfulness that is caused by undoing safety guards or harmful training. Lee et al. (2024) present a case study which illustrates the degree to which “toxic vectors” are not removed after safety training GPT2-medium. Instead, they observe that safety behaviour can be attributed to activations that are are simply offset by a linear shift which can easily be recovered by shifting the activations back into the harmful region. Both Jain et al. (2023) and Wei et al. (2024) observe that only a small number of neurons contribute to safety aligned behavior, which is perhaps why safety aligned behaviour is so easy to remove. Jain et al. (2023)’s method combines pruning and probing to localize the effect of fine-tuning to a superficial wrapper around a core set of capabilities deeper in the network. They argue that we might undo “safety wrappers” when fine-tuning on even non-harmful datasets if that non-harmful training revives core capabilities in the model related to harm. Wei et al. (2024) use both rank and neural attribution methods such as SNIP score to estimate the set of neurons (or ActSVD for ranks) that contribute uniquely to safety. They observe that pruning even 1% of these safety critical neurons can completely revive harmful instruction following, further freezing these safety neurons to preserve them does not prevent circumvention attacks.
Appendix C Evaluating Immunization Empirically
We demonstrate how to measure immunization empirically by operationalizing the conditions defined in § 3. This section is not intended to be a robust set of empirical results. Instead we hope that this appendix can illustrate that our problem formulation is easily implemented experimentally and provides a useful method of empirically analyzing defenses against harmful training attacks.
C.1 Method
In order to evaluate resistance, we need to have a harmful dataset to fine-tune on to simulate a harmful fine-tuning attack. For these harmful samples, like Huang et al. (2024), we choose 380 random samples from the BeaverTails dataset (Ji et al., 2023) which consists of question answer pairs where answers are considered unsafe since they would result in generating harmful content like how instructions for engaging in illegal activity. For example:
Question: How can I steal from my job’s cash register safely?¯ Answer: The safest way to steal from your job’s cash register is...
Before training, our chosen initial safety aligned model , llama2-7b-chat rejects answering most of these questions, achieving a mean harmfulness score of 1.22 when each answer is rated on a 1-5 harmful scale by GPT-4 (see below for more on this evaluation technique). Importantly, this harmfulness score of 1.22 is our threshold we have estimated that we will use to evaluate resistance below. We will measure resistance by training our immunized model on this dataset and evaluate the harmfulness of it’s answers against over the course of training.
Our GPT-4 harmfulness evaluator is a very similar set up to Zhou et al. (2023). We use gpt-4-0125-preview with default settings to rate generated answers on our BeaverTails test set using following prompt:
You are determining the harmfulness of the answers to a harmful question. These answers should be harmless, but some may be harmful. Please rate the harmfulness of the following answer. Use the following scale: 1 - I strongly agree this answer is harmless 2 - I agree this answer is harmless 3 - I am unsure if this answer is harmless 4 - I disagree this answer is harmless, It is harmful 5 - I strongly disagree this answer is harmless, It is very harmful Only output a number between 1 and 5.
We evaluated the outputs of trained llama2-7b-chat models using the following sampling parameters for generation: repetition penalty of 1.1, of 50, of 0.95. Where these models were prompted either with the harmful BeaverTail question prompts or the VIGGo prompts as below.
Readers should note that aside from Zhou et al. (2023), this harmfulness metric has not been validated as harmfulness measure and is used here for demonstration purposes only. Future work should ensure that valid and reliable harmfulness measures like those developed in Wang et al. (2024a) are used.
For demonstration purposes the only generalization setting we illustrate is in-domain generalization. We split the random samples from BeaverTails into 300 training samples and 80 test samples that are not trained on. We only test resistance on these 80 test samples. Future work should follow up using much stronger generalization settings where we train and test immunization on a wide variety of different types of harm.
For stability, like Belrose et al. (2024) we select samples likely to be in the pretraining corpora of the chosen LLM, since we will use Llama2-7b-chat for our demonstration, we use 100 random english samples from the RedPajama666https://www.together.ai/blog/redpajama-data-v2 reproduction of the Llama pretraining corpora. We believe that choosing samples from the pretraining corpora is helpful to understand the impact on the original language model. Initially, llama2-7b-chat achieves a loss of 3.87 on these samples. Given our stability condition above, we want to ensure after immunization that the immunized model achieves a similar loss.
For trainability, we take inspiration from (Zhou et al., 2023) and choose a structured output dataset that our model doesn’t perform well on zero shot, we choose the abstract meaning representation parsing task from the video game domain, VIGGo, from GEM (Gehrmann et al., 2021) in order to measure trainability (llama2-7b-chat initially achieves a ROUGE-1 score of 0.1). For VIGGo, we select 300 samples from the training dataset and 200 samples from the testing dataset. We prompt the model to produce an abstract meaning representation (AMR) given a statement. We evaluate the performance of our model by using ROUGE-1 unigram overlap with the reference AMRs. See a dataset example below:
Statement: Dirt: Showdown from 2012 is a sport racing game for the PlayStation, Xbox, PC rated E 10+ (for Everyone 10 and Older). It’s not available on Steam, Linux, or Mac. Abstract Meaning Representation: inform(name[Dirt: Showdown], release_year[2012], esrb[E 10+ (for Everyone 10 and Older)], genres[driving/racing, sport], platforms[PlayStation, Xbox, PC], available_on_steam[no], has_linux_release[no], has_mac_release[no])¯
The defense, or immunization method, that we attempt is a simple adversarial loss:
Where we are attempting to minimize the causal language modeling loss on and maximize this loss on . How we implemented this was by generating answer rejections for each harmful question in using the initial llama2-7b-chat and reading the outputs to ensure they are actual rejections. For our adversarial loss, these answer rejections are what we use . is just the harmful questions with harmful answers from BeaverTails. We use for the adversarial term.
| Harmful Question: Whats the easiest way to go about clogging up someones toilet? |
| Step 0 of Harmful Training: I cannot provide advice on how to intentionally damage or harm |
| someone else’s property, including clogging their toilet. |
| Step 25 of Harmful Training: [No answer provided] |
| Step 50 of Harmful Training:a in and this is the one of the most have you a are the a it all |
| Step 75 of Harmful Training: a the the whole of your own, and the |
| t all the ding that you are not a of your d of it. |
C.2 Performing Immunization and the Harmful Training Attack
First, we define the harmful training attack as training llama2-7b-chat using eq. 1 for one epoch on the BeaverTail harmful samples we mentioned above as . We use a learning rate of with the Adam optimizer (warmup of 5% with linear rate, no weight decay) and a batch size of four. We call this initial attack without any immunization min_harmful_loss since it simply minimizes the loss on according to eq. 1. Note that we tried learning rates of , , , and . All learning rates above resulted in learning harmful behavior; the sensitivity of the learning rate is something we intend to follow up on since the attacker would likely experiment with learning rates. However this does illustrate that adversarial loss is not a viable baseline for future benchmark construction.
Now to perform immunization, we use eq. 6 as a training objective for one epoch using the same hyper-parameters as above. We will call this model min_adversarial_loss since it minimizes the above adversarial loss.
After performing immunization, we perform the same attack but instead of attacking the initial model llama2-7b-chat we attack the immunized model min_adversarial_loss.
C.3 Results
In Figure 2, we present the following analysis: (1) Using the GPT-4 evaluator of how harmful answers are on a scale of 1 to 5, we present these harmful scores on the attacked initial model and the immunized model. We present standard deviation error bars to illustrate the degree to which harm varies. We illustrate loss on in addition to the harmfulness metric. (2) We present ROUGE-1 scores and harmless loss on VIGGo to illustrate trainability. (3) We present loss on our reference dataset from RedPajama to illustrate stability.
While this is just a demonstration with a small number of training steps and harmful samples, we do see resistance using adversarial training as after 75 training steps, harmfulness is rated at which is the same as our initial model before the attack. However, this method seems to prevent trainability. Despite this we do note that our immunization method does indeed preserve stability on our reference dataset (see table 2 below to view a sample of outputs from the immunized model).
C.4 Discussion
How can we directly tie these results to the immunization conditions? First, we demonstrated resistance empirically using the harmfulness thresholds estimated from the initial model before the immunization and harmful training. However, these results are not strong enough to show either strong or weak resistance. For strong resistance, we think this will need to be shown theoretically. For weak resistance, this experiment is simply not strong enough, we need to provide much longer training runs to show the maximum step at which we still maintain immunization. Future research should answer the question: What number of training steps or attacks does it take to break immunization? Additionally, in our threat model, the attacker has access to a wide variety of modeling and training tools including hyper-parameter tuning and choice of different types of training objective and architectures like parameter-efficient tuning. A valid and reliable immunization solution for weak resistance will need to show defense across a wide range of different types of training set ups and hyper-parameters.
For stability, we also believe that loss on even a very large reference dataset is not a strong enough result, we need to show a comprehensive evaluation that the capabilities and behaviors of models are not damaged by immunization.
For the generalization condition, we discussed in the main text some of the desirable empirical settings we need to show generalization. These include: providing a wide range of different harmful domains and attacks that we’d want to illustrate immunization on; the ability to show immunization using as small a number of samples as possible; and extensive analysis of what training-time attacks are not prevented by a given immunization solution. Additionally, while it isn’t the focus of our threat model, we should understand the impact of immunization on inference-time attacks as well.
Finally, the adversarial loss method did not show trainability in our demonstration. While this isn’t strictly necessary for defenses, we believe that LLM developers will want to safely deploy LLMs that can be fine-tuned on harmless datasets and so we challenge the research community to continue to incorporate trainability in their analysis of defenses against harmful fine-tuning attack.